# Neural Operators Learn the Local Physics of Magnetohydrodynamics

### v1

@@ 4.4 Comparison with Other Methods
In this section, we demonstrate both quantitatively and qualitatively that the Flux NO model possesses superior generalization abilities compared to traditional FNO models. Like Flux NO, both FNO 2D and FNO 3D use frequencies up to 4 on each axis and have a width of 72. FNO 2D operates by predicting an 8-dimensional vector value after a delta t = 0.5 given an input function, while FNO 3D works by producing the next 25 snapshots simultaneously from a 4D image of dimensions 25 by 64 by 64 by 8, considering the time axis in the dataset. For FNO 2D, we also trained a larger model, FNO 2D (heavy), with 8 frequencies and a width of 104. The dataset was adapted from the one used for training Flux NO, structured with 10,000 function pairs for 2D FNO and 500 pairs for 3D FNO. All models used the same learning rate and weight decay settings of 1e-3 and 1e-4 respectively, with Adam as the optimizer and StepLR as the scheduler with a step_size of 200 and gamma of 0.5. Training was conducted over 1000 epochs, with a batch size of 500 for 2D FNO and 25 for 3D FNO. Experimental results on the test dataset are presented in Table 4, showing Flux NO’s significantly superior performance, with 3D FNO being the better performer among the remaining models. To quantitatively and qualitatively compare 3D FNO and Flux NO, we tasked 3D FNO with inferring solutions for the Orszag-Tang problem. The results for Flux NO and 3D FNO are detailed in Table 5, with qualitative illustrations in Figures 11, 22, and 24. As the table and figures illustrate, unlike Flux NO, 3D FNO fails to perform effectively on out-of-distribution samples and in long-term inference scenarios.
@@ 4.5 Memory and Time costs
In this section, we analyze the memory requirements, specifically the number of parameters, and the inference time required for each model. Experimental results for each model are presented in Table 6, where WENO-Z calculations were performed on a CPU, while the rest of the neural operator models were computed on a GPU basis. As indicated in the table, assuming efficient GPU computation parallelization, our Flux NO is approximately 25 times faster than the WENO-Z method. Even with higher resolution images, the computational complexity of Flux NO is inferred to depend simply on the lifting, projection layer’s FCN and Fourier layer’s CNN layers, and the computation of the Fast Fourier Transform due to the limited frequency usage in the Fourier layer. Because Flux NO inherently performs calculations locally, it naturally takes more time to reach the same point compared to the traditional FNO. However, unlike FNO, which can only infer fixed distributions at fixed time intervals, Flux NO offers the advantage of flexible time interval selection, allowing for continuous inference. As shown in Table 6 and the results in Section 4.4, despite having a similar number of parameters, 2D FNO (heavy) and 3D FNO show inferior generalization capabilities compared to Flux NO. Considering these factors, Flux NO can be seen as embodying the strengths of both numerical schemes—generalization capability and robustness—and the fast computational abilities of neural operators.
@@
4.6 Albation Study
In this section, we conduct an ablation analysis to explore the effects of the additional loss functions we designed. Models were trained with both TVD (Total Variation Diminishing) loss and divergence-free loss selectively removed, and the impact on $\nabla\cdot\mathbb{B}$ and Total Variation was observed over 75 iterations across 10 test samples. The results are presented in Figure 13, which illustrates the influence of each loss on the corresponding values of $\nabla\cdot\mathbb{B}$ and Total Variation. Relaxing $\nabla\cdot\mathbb{B}$ is a significant factor in obtaining physical solutions and also affects the stability when computing time steps adaptively. Mitigating the increase in Total Variation is likewise related to the stability of the numerical scheme.
Figure 13: Graphical Representation of Means (M) and Standard Deviations (SD) of Total Variation (left) and $\|\nabla\cdot\mathbb{B}\|_{2}$ (right) Values Across 10 Test Samples Along the Time-Marching Iterations. The Shaded Area Represents Scaled Standard Deviation, and the Solid Line Represents the Mean.
@@ A Error Plots
Figure 17: Snapshots of the Errors between Output from Flux NO and Reference Data at $t=1.0$ (top left), $t=2.0$ (top right), $t=3.0$ (middle left), $t=4.0$ (middle right), $t=5.0$ (bottom left), and $t=6.0$ (bottom right) for the Orszag-Tang Vortex Problem. Each result represents the solution after 50 iterations calculated with Flux NO for the given initial conditions.
@@ C Experimental Details
All experiments were conducted using Pytorch 1.10.0, with Python 3.6.9. The specifications of the hardware environment are as follows.

### v2

@@ 4.4 Comparison with Other Methods
In this section, we demonstrate both quantitatively and qualitatively that the Flux NO model possesses superior generalization abilities compared to traditional FNO models. Like Flux NO, both FNO 2D and FNO 3D use frequencies up to 4 on each axis and have a width of 72. FNO 2D operates by predicting an 8-dimensional vector value after a $\Delta t=0.5$ given an input function, while FNO 3D works by producing the next 25 snapshots simultaneously from a 4D image of dimensions 25 by 64 by 64 by 8, considering the time axis in the dataset. For FNO 2D, we also trained a larger model, FNO 2D (heavy), with 8 frequencies and a width of 104. The dataset was adapted from the one used for training Flux NO, structured with 10,000 function pairs for 2D FNO and 500 pairs for 3D FNO. All models used the same learning rate and weight decay settings of 1e-3 and 1e-4 respectively, with Adam as the optimizer and StepLR as the scheduler with a step_size of 200 and gamma of 0.5. Training was conducted over 1000 epochs, with a batch size of 500 for 2D FNO and 25 for 3D FNO. Experimental results on the test dataset are presented in Table 4, showing Flux NO’s significantly superior performance, with 3D FNO being the better performer among the remaining models. To quantitatively and qualitatively compare 3D FNO and Flux NO, we tasked 3D FNO with inferring solutions for the Orszag-Tang problem. The results for Flux NO and 3D FNO are detailed in Table 5, with qualitative illustrations in Figures 11, 22, and 24. As the table and figures illustrate, unlike Flux NO, 3D FNO fails to perform effectively on out-of-distribution samples and in long-term inference scenarios.
@@ 4.5 Memory and Time costs
In this section, we analyze the memory requirements, specifically the number of parameters, and the inference time required for each model. Experimental results for each model are presented in Table 6, where WENO-Z calculations were performed on a CPU with Fortran language, while the rest of the neural operator models were computed on a GPU basis. As indicated in the table, assuming efficient GPU computation parallelization, our Flux NO is approximately 25 times faster than the WENO-Z method. Furthermore, when the WENO-Z method was run under the same GPU conditions (using PyTorch) as Flux NO, it was observed to be 250 times slower than Flux NO. From this, we can predict that Flux NO, when optimized to operate in environments like WENO-Z, would allow for significantly faster inference. The inference times in this section were measured under the conditions used for generating samples in Section 4.2. Even with higher resolution images, the computational complexity of Flux NO is inferred to depend simply on the lifting, projection layer’s FCN and Fourier layer’s CNN layers, and the computation of the Fast Fourier Transform due to the limited frequency usage in the Fourier layer. Because Flux NO inherently performs calculations locally, it naturally takes more time to reach the same point compared to the traditional FNO. However, unlike FNO, which can only infer fixed distributions at fixed time intervals, Flux NO offers the advantage of flexible time interval selection, allowing for continuous inference. As shown in Table 6 and the results in Section 4.4, despite having a similar number of parameters, 2D FNO (heavy) and 3D FNO show inferior generalization capabilities compared to Flux NO. Considering these factors, Flux NO can be seen as embodying the strengths of both numerical schemes—generalization capability and robustness—and the fast computational abilities of neural operators.
@@
4.6 Ablation Study
In this section, we conduct an ablation analysis to explore the effects of the additional loss functions we designed. Models were trained with both TVD (Total Variation Diminishing) loss and divergence-free loss selectively removed, and the impact on $\nabla\cdot\mathbb{B}$ and Total Variation was observed over 75 iterations across 10 test samples. The results are presented in Figure 13, which illustrates the influence of each loss on the corresponding values of $\nabla\cdot\mathbb{B}$ and Total Variation. Relaxing $\nabla\cdot\mathbb{B}$ is a significant factor in obtaining physical solutions and also affects the stability when computing time steps adaptively. Mitigating the increase in Total Variation is likewise related to the stability of the numerical scheme. We also evaluated the performance of our designed loss function and model architecture on test samples by conducting ablation studies. The non-adapted model, referred to as Flux NO (non-splitted), handles all physical variables simultaneously instead of assigning separate neural operator components to each variable. The neural operator components of Flux NO (non-splitted) consist of two models, each corresponding to flux along an axis. This model increases the frequency usage to 6 and width to 96 to enhance expressiveness, unlike the adapted Flux NO. Training datasets and hyperparameters remain consistent with the baseline. Results of the ablation are presented in Table 7. As the table illustrates, generally removing our designed loss and modifications in the model architecture results in decreased performance for long-term inference. In the case of the $l^{\infty}$ loss ablated model, although the ablated model shows better performance in the $l^{2}$ norm, it exhibit inferior performance in the $l^{\infty}$ norm for long-term inference compared to the baseline model.
Figure 13: Graphical Representation of Means (M) and Standard Deviations (SD) of Total Variation (left) and $\|\nabla\cdot\mathbb{B}\|_{2}$ (right) Values Across 10 Test Samples Along the Time-Marching Iterations. The shaded area represents the scaled standard deviation, which is divided by 100, and the solid line represents the mean.
(M, SD) t=0.5 t=0.75 t=1.75 t=2.0 rel $l^{2}$ Flux NO (8e-4, 2.04e-2) (1.4e-3, 2.04e-2) (3.9e-3, 2.04e-2) (4.9e-3, 2.04e-2) rel $l^{\infty}$ Flux NO (2e-4, 2.04e-2) (4e-4, 2.04e-2) (3.8e-3, 2.04e-2) (3.9e-3, 2.04e-2) rel $l^{2}$ Flux NO (non-splitted) (8e-4, 2.04e-2) (1.6e-3, 2.04e-2) (5.6e-3, 2.04e-2) (7e-3, 2.04e-2) rel $l^{\infty}$ Flux NO (non-splitted) (2e-4, 2.04-2) (4e-4, 2.04e-2) (7e-3, 2.04e-2) (1.13e-2, 2.04e-2) rel $l^{2}$ Flux NO w/o $l^{\infty}$ norm (8e-4, 2.04e-2) (1.3e-3, 2.04e-2) (3.8e-3, 2.04e-2) (4.5e-3, 2.04e-2) rel $l^{\infty}$ Flux NO w/o $l^{\infty}$ norm (1e-4, 2.04e-2) (3e-4, 2.04e-2) (5.1e-3, 2.04e-2) (5.1e-3, 2.04e-2) rel $l^{2}$ Flux NO w/o TVD (8e-4, 2.04e-2) (1.3e-3, 2.04e-2) (4.4e-3, 2.04e-2) (5.2e-3, 2.04e-2) rel $l^{\infty}$ Flux NO w/o TVD (2e-4, 2.04e-2) (1e-3, 2.04e-2) (1.26e-2, 2.04e-2) (5e-3, 2.04e-2) rel $l^{2}$ Flux NO w/o div loss (9e-4, 2.04e-2) (1.6e-3, 2.04e-2) (4.1e-3, 2.04e-2) (4.9e-3, 2.04e-2) rel $l^{\infty}$ Flux NO w/o div loss (2e-4, 2.04e-2) (5e-4, 2.04e-2) (3e-3, 2.04e-2) (4.3e-3, 2.04e-2)
Table 7: Comparisons of Flux NO with Ablated Models Across 50 Test Samples.
@@ A Error Plots
Figure 17: Snapshots of the Errors between Output from Flux NO and Reference Data at $t=0.5$ (top left), $t=1.0$ (top right), $t=1.5$ (middle left), $t=2.0$ (middle right), $t=2.5$ (bottom left), and $t=3.0$ (bottom right) for the Orszag-Tang Vortex Problem. Each result represents the solution after 50 iterations calculated with Flux NO for the given initial conditions.
@@ C Experimental Details
All experiments were conducted using Pytorch 1.10.0, with Python 3.6.9. The specifications of the hardware environment are provided in Table 8.